November 2014 Reports

To the main

27 Nov 2014 (Thursday)

  • MC and user jobs. Current schedule for "Legacy Run1 stripping campaign": Davinci release is out today so validation should start very soon. Still hoping for full campaign next Monday if all well.
  • T0: NTR
  • T1: NTR
  • Other: Problem with the SAM dashboard no showing site availability properly overnight on Tuesday/Wednesday

24 Nov 2014 (Monday)

  • MC and user jobs. Current schedule for "Legacy Run1 stripping campaign": Validate on a few K files end of week. Start full campaign next Monday if all well.
  • T0: NTR
  • T1: NTR

20 Nov 2014 (Thursday)

  • MC and user jobs. Staging files to disk buffer for Legacy Run1 stripping campaign has been restarted
  • T0: lbvobox18 instabilities yet to be understood, likely TCP flooding
  • T1: NTR

17 Nov 2014 (Monday)

  • MC and user jobs. "Legacy Run1 stripping campaign", new schedule for stripping21 is on Wed (the earliest)
  • T0: NTR
  • T1: Replication of full.dst is over in all sites except GridKa due to slow staging progress

13 Nov 2014 (Thursday)

  • MC and user jobs. "Legacy Run1 stripping campaign" first batch of jobs currently validated by physicists, continue to run early next week
  • T0: One VOMS server (old) crashed and needed to be restarted (GGUS:110068)
  • T1: glitch with FTS transfers to RRCKI on Tuesday, now fixed

10 Nov 2014 (Monday)

  • MC and user jobs. "Legacy Run1 stripping campaign" to be launched today @ T1 sites
  • T0: NTR
  • T1: problem with SLC6.6 worker nodes and ROOT 6 based applications is fixed by deployment of a new software stack. All new productions (MC) launched as of today will be capable of running also on SLC6.6 nodes (some T1 sites were affected)

3 Nov 2014 (Monday)

  • MC and User jobs. Prestaging from tapes for Stripping21 campaign.
  • T0: Problems accessing files at CERN over the weekend (GGUS:109808 and GGUS:109672). Affected ~2K files. Problem fixed fast and jobs and transfers have finished successfully.
  • T1:
    • IN2P3 : dCache upgrade to fix root6 access problems.
    • GridKa : Investigating possible problems accessing two files and also some FTS transfer issues (GGUS:109825).
  • Others : Changes in way root6 handles include files causing crashes at many sites on grid. Working on a fix within LHCb software stack.

-- JoelClosier - 31 Mar 2014

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2015-01-05 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback